a) Compute the takt time for a system where the total time per shift is 480 minutes, there are two shifts, and workers on each shift are given two 15-minute breaks and 45 minutes for lunch. Daily demand is 650 units.

b) What cycle time would match capacity and demand if demand is 120 units a day, there are two shifts of 480 minutes each, and workers are given two half-hour breaks during each shift, one of which is for lunch or dinner?

c) Compute the takt time for a service system that has a total work time of 440 minutes per day, two 10-minute breaks, and an hour for lunch. The service system must process 100 jobs per day.
Answer #1

A. Net time per shift = 480 - (2 * 15) - 45 = 405

Total net time = 405 * 2 = 810

Takt time = Total net time / Daily demand = 810 / 650 = 1.246 minutes

B. Net time per shift = 480 - (2 * 30) = 420

Total net time = 2 * 420 = 840

Daily demand = 120

Cycle time = Total net time / Daily demand = 840 / 120 = 7 minutes

C. Net time per shift = 440 - (2 * 10) - 60 = 360

Daily demand = 100

Takt time = 360 / 100 = 3.6 minutes.